# references.bib

### v2

@@ Abstract.
Algorithmic discrimination is a critical concern as machine learning models are used in high-stakes decision-making in legally protected contexts. Although substantial research on algorithmic bias and discrimination has led to the development of fairness metrics, several critical legal issues remain unaddressed in practice. To address these gaps, we introduce a novel decision-theoretic framework grounded in anti-discrimination law of the United Kingdom, which has global influence and aligns more closely with European and Commonwealth legal systems. We propose the “conditional estimation parity” metric, which accounts for estimation error and the underlying data-generating process, aligning with legal standards. Through a real-world example based on an algorithmic credit discrimination case, we demonstrate the practical application of our formalism and provide insights for aligning fairness metrics with legal principles. Our approach bridges the divide between machine learning fairness metrics and anti-discrimination law, offering a legally grounded framework for developing non-discriminatory automated decision systems.
UK anti-discrimination law, anti-discrimination doctrine, machine learning, automated decision making, algorithmic discrimination, algorithmic fairness, estimation error, true data generating process
copyright: acmlicensedjournalyear: 2025doi: XXXXXXXconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; TBA, 2025; Athens, Greece
@@ 1. Introduction
Instances of large-scale failures of supervised machine learning (SML) based decisions, from disproportionately harming vulnerable people in algorithmic immigration assessments (forster2022; oxkul2023), unfair welfare eligibility assessments (eubanks2018; senecat2023; Lighthouse2024), to racial bias in criminal justice systems (zilka2022; adams2024), have spurred a rich literature on algorithmic bias and discrimination. Despite such literature from the ML and legal communities focusing on fairness metrics (mayson2019; ho2020; xiang2021; barocas2016; hacker2018; wachter20212; xenidis2020; hellman20201; gillis2022; adams2022; weerts2023), several critical legal issues remain unaddressed in practice. Critical advice for modellers on avoiding and mitigating unlawful discrimination in different jurisdictions remains limited. This paper prioritises three primary issues within current approaches that, if ignored, limit both the legal and practical effectiveness of fairness interventions.
@@
First, while algorithmic fairness research largely identifies statistical disparities in predicted outcomes for binary, marginalised groups; legally, discrimination is both broader and more detailed. The conflation of fairness and non-discrimination has led to a predominant focus on outcome disparity as the main indicator of bias, yet in legal contexts, discrimination entails a more layered and multidimensional analysis.
Second, there is often a disconnect between technical modelling practices and the requirements of anti-discrimination law. Many technical approaches overlook estimation error, fail to account for the underlying data-generating process (DGP) or assumes the true causal structure is known. Yet these have significant legal and practical consequences.
Third, the predominance of analysis of fairness and discrimination in ML from the United States (US), lack of non-US ML examples (quy2022), and limited legal scholarship translating these concepts across other jurisdictions, has inadvertently fostered a series of misconceptions on what would be unlawful algorithmic discrimination.
@@
Few papers have engaged with anti-discrimination laws outside the US (see e.g., xenidis2020; lattimore2020; wachter20211; weerts2023; kiesow2023; langenbucher2023), and fewer still in the United Kingdom (UK(see e.g., sanchez2020; adams2022; kelly2021). By avoiding the nuanced legal realities of other jurisdictions, models designed to comply with US laws may breach UK laws or those in comparable jurisdictions. English common law is either in force or is the dominant influence in 80 legal systems that govern approximately 2.8 billion people, not including the US (CIA). In particular, UK anti-discrimination law is very similar to numerous Commonwealth and common law jurisdictions, including Australia (AHRC2014), Canada (CanadaHRA), India (IndiaConst), New Zealand (NZHRA), South Africa (SADL), and the pending bill in Bangladesh (BADB2022). European Union (EU) law also has broadly the same principles and discrimination case law evolved in parallel during the UK’s membership (EUCFR).
@@
As new AI regulations emerge worldwide to prevent discriminatory practices, the need for a more jurisdictionally nuanced understanding of unlawful discrimination decision-making based on SML is imperative (EUAIAct; EO14110). Our paper addresses this gap by providing a rigorous analysis of UK anti-discrimination law, correcting certain mischaracterisations, and establishing a more accurate foundation for developing non-discriminatory ML in the UK and related jurisdictions.
@@
    We propose conditional estimation parity as a new, legally informed target to minimise the legal and practical effects of estimation error in SML models.
@@ 2.1. Automated Decision-Making
Let $x_{i}\in\mathbb{R}^{p}$ be a vector of observed attributes for individual $i$. A decision-maker must choose a decision $a\in\mathcal{A}$, where $\mathcal{A}$ is closed. We assume the decision-maker wants to decide based on a future outcome $y_{i}\in\mathcal{Y}$ for individual $i$. Further, we assume $\mathcal{Y}=\mathbb{N}$, which can be relaxed. Decision-making under uncertainty has long been studied in statistical decision theory (savage1956; degroot1970optimal; berger1985statistical; parmigiani2010). In a decision setting, let $u(y,a)$ be a utility function that summarises the utility for the decision-maker, where the optimal decision under uncertainty is
@@ 2.2. Algorithmic Fairness
Algorithmic fairness metrics generally measure prediction disparities across groups with different legally protected characteristics commonly identified in datasets, including gender and race (hardt2016; corbett2017; narayanan2018; verma2018; mulligan2019; binns2021; mehrabi2021; mitchell2021). This research has resulted in several proposals, including statistical metrics to assess the fairness of individual predictive models (verma2018; speicher2018; carey2023; caton2023), fairness for model auditing (kim2017; raji2020; kasy2021; mokander2023; oneil2023), and fairness constraints on models (corbett2017; zafar2019; Heidari2019; Xu2020; becker2023). We outline two core metrics that are relevant to our work – statistical parity and conditional statistical parity. To define these metrics in our notation, we separate $x_{i}$ into protected and legitimate features $x_{i}=(x_{pi},x_{li})$; we drop $i$ to simplify notation. Here, $x_{p}\in\mathcal{C}$ indicates protected attributes, with $\mathcal{C}$ being the set of different categories or groups.
Statistical parity, or demographic parity, is one of the central algorithmic fairness metrics (corbett2017; verma2018; narayanan2018; mehrabi2021). For statistical parity to hold, it requires that
@@
such that the model predictions, in expectation over $x$, need to be the same for the different groups (corbett2017; verma2018). Given that the decision function $d(\boldsymbol{\pi})$ is the same for the different groups, statistical parity results in equal decisions across those groups.
@@
so there should be no difference in model predictions or decisions between groups given by the protected attribute, conditional on legitimate features $x_{l}$ (corbett2017; verma2018; castelnovo2022).
Other related group comparison metrics include error parity, balanced classification rates, and equalised odds (hardt2016; corbett2017; verma2018; narayanan2018; mehrabi2021; corbett2023). Individual approaches to parity have also considered whether otherwise identical individuals are treated differently if they have different protected attributes (dwork2012; kilbertus2018; binns2020). Concepts from causal inference and counterfactual reasoning have been proposed to measure outcome consistency for individuals across protected groups (kusner2017; kilbertus2017; russell2017; zhang2018fairness; chiappa2019; wu2019; nilforoshan2022causal; alvarez2023counterfactual). It is not in the scope or aim of this paper to evaluate these numerous algorithmic fairness metrics.
@@ 2.3. Anti-discrimination Law and Decision-Theoretic Approach
Identifying a discriminatory AI system has occurred in a variety of contexts, often without identifying whether it is unlawfully discriminatory. A non-lawyer may be surprised by several types of unfair behaviours that are not legally prohibited and some inconspicuous decisions that result in unlawful discrimination (khaitan2015). As a term that has entered the common vernacular, despite its specific meanings in several contexts, it has led to the conflation of discrimination and unlawful discrimination. Importantly, anti-discrimination law only applies to a select group of duty-bearers (khaitan2015). For example, an individual choosing not to be friends with people based on their sexuality or choosing not to marry someone based on their race is not unlawful (emens2009; khaitan2015). Despite certain actions appearing unreasonable or unfair, they are not always prohibited. Consider an algorithm that rejects a loan application because the applicant uses an Android phone rather than an iOS device (aggarwal2021; langenbucher2023). While it may show a correlation to the applicant’s income, one may agree that it is unfair because it does not reflect the individual likelihood of defaulting on a loan and instead penalises the individual for how much money they spent on a mobile device. However, under UK law, it would not be unlawful because spending habits, income, or even poverty, are not protected attributes (poverty2022). Therefore, to engage in a functional analysis of UK anti-discrimination laws in automated decision systems, it is helpful to frame its legal elements.
@@
In most algorithmic fairness literature, the primary, or often the sole, consideration is whether the ML outputs result in disparate predictions or decisions for a protected group (hardt2016; verma2018; narayanan2018). Such approach only considers aspects of the rule articulation function of anti-discrimination law. However, we take a wholistic, legally-informed approach:
@@
We formalise UK anti-discrimination law in a decision-theoretic framework that provides a systematic approach for the decision-maker to make optimal ML design choices under uncertainty and under legal constraints. Our decision-theoretic view, where actions map states to outcomes, each with associated utilities, enabling a formal assessment of modelling choices within UK anti-discrimination law.
@@ 3. What types of discrimination does the law prohibit?
Unlawful and prohibited discrimination is defined by (1) protected contexts, (2) protected characteristics, and (3) prohibited conduct. Protected contexts are defined by the imposed duties on government, employers, landlords, providers of goods and services (EqualityAct). The protected characteristics under the Equality Act are age, disability, gender reassignment, marriage and civil partnership, pregnancy and maternity, race, religion or belief, sex and sexual orientation (EqualityAct, s 4). Two types of conduct are prohibited: direct discrimination and indirect discrimination.
@@ 3.1. Algorithmic Direct Discrimination
Less favourable treatment of an individual in a protected context based on one or multiple protected characteristics is unlawful direct discrimination (EqualityAct, s 13(1), 14). Where a model $\hat{p}(y|x)$ uses a protected attribute $x_{p}$, and there is a difference in predictions between the protected groups defined by $x_{p}$ that results in less favourable treatment for an individual in that group, this risk arises.
[…]
@@
Further, from a technical perspective simply removing protected characteristics may reduce accuracy and utility (Zhang2017; khani2021), and does not remove the risk of bias or disparity (dwork2012; lipton2018; kleinberg2019). If the inclusion of $x_{p}$ or $\tilde{x}_{p}$ improves the model accuracy without resulting in less favourable treatment for protected individuals, it may avoid direct discrimination. There is an absence of any legal guidance in the UK on the use of protected attributes in automated decision-making. [^1] Although in the context of data processing, it may be lawful to use data on personal characteristics if it is necessary to identify or review equality of opportunity or treatment between groups of people, or to prevent or detect other unlawful acts (DPA, Sch 1). Pending further legal guidance, it is important to carefully consider the effects of including $x_{p}$ or $\tilde{x}_{p}$.
@@
Second, intention is irrelevant to direct discrimination, “no hostile or malicious motive is required” (Essop; lords1990; EandJFS). Importantly, this diverges from US law and highlights that intention is immaterial to UK direct discrimination and to consideration of intentional proxy discrimination (cf. prince2020; tschantz2022).

### v3

# Formalising Anti-Discrimination Law in Automated Decision Systems

@@ Abstract.
Algorithmic discrimination is a critical concern as machine learning models are used in high-stakes decision-making in legally protected contexts. Although substantial research on algorithmic bias and discrimination has led to the development of fairness metrics, several critical legal issues remain unaddressed in practice. The paper addresses three key shortcomings in prevailing ML fairness paradigms: (1) the narrow reliance on prediction or outcome disparity as evidence for discrimination, (2) the lack of nuanced evaluation of estimation error and assumptions that the true causal structure and data-generating process are known, and (3) the overwhelming dominance of US-based analyses which has inadvertently fostered some misconceptions regarding lawful modelling practices in other jurisdictions. To address these gaps, we introduce a novel decision-theoretic framework grounded in anti-discrimination law of the United Kingdom, which has global influence and aligns closely with European and Commonwealth legal systems. We propose the “conditional estimation parity” metric, which accounts for estimation error and the underlying data-generating process, aligning with UK legal standards. We apply our formalism to a real-world algorithmic discrimination case, demonstrating how technical and legal reasoning can be aligned to detect and mitigate unlawful discrimination. Our contributions offer actionable, legally-grounded guidance for ML practitioners, policymakers, and legal scholars seeking to develop non-discriminatory automated decision systems that are legally robust.
UK anti-discrimination law, algorithmic discrimination, algorithmic fairness, machine learning, statistical decision-theory, estimation error, epistemic uncertainty
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732015isbn: 979-8-4007-1482-5/2025/06conference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greececcs: Applied computing Lawccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
@@ 1. Introduction
Instances of large-scale failures of supervised machine learning (SML) based decision systems, from disproportionately harming vulnerable people in algorithmic immigration assessments (Forster, 2022; Ozkul, 2023), unfair welfare eligibility assessments (Eubanks, 2018; Sénécat, 2023; Geiger et al., 2024), to perpetuating racial biases within criminal justice systems (Zilka et al., 2022; Adams-Bhatti and Sargeant, 2024), have spurred a rich literature on algorithmic bias and discrimination. Despite such literature from the ML and legal communities focusing on fairness metrics intended to identify and address algorithmic bias (Mayson, 2019; Ho and Xiang, 2020; Xiang, 2021; Barocas and Selbst, 2016; Hacker, 2018; Wachter et al., 2021a; Xenidis, 2020; Hellman, 2020; Gillis, 2022; Adams-Prassl et al., 2022; Weerts et al., 2023), several critical legal issues remain unaddressed in practice. Practitioners, especially model developers and decision-makers, continue to lack precise guidance on effectively avoiding and mitigating unlawful algorithmic discrimination in compliance with anti-discrimination laws across various jurisdictions. This paper explicitly identifies and prioritises three primary shortcomings prevalent in current approaches that, if ignored, undermine both the lawfulness and practical effectiveness of fairness interventions.
@@
First, while algorithmic fairness research largely emphasises discrimination as statistical disparities in predicted outcomes for binary, marginalised groups. The common conflation of fairness and legal non-discrimination has led to a predominant focus on outcome disparity as the main indicator of unfairness, yet unlawful discrimination is both broader and more detailed. This gap creates a critical mismatch between what ML methods identify as unfair and what anti-discrimination law recognises as unlawful, rendering many fairness interventions inadequate, counterproductive, or even unlawful.
Second, a persistent disconnect exists between technical modelling practices and the requirements of anti-discrimination law, resulting in legal blind spots and methodological oversights. Conventional technical approaches frequently overlook important aspects of statistical uncertainty and estimation error, fail to account for the underlying data-generating process (DGP), or assume that the true causal structure behind discrimination is already known. In reality, the outputs of SML models are approximations, subject to inherit sampling variability, label noise, data limitations, and model misspecification. These methodological issues are not academic abstractions – they carry concrete legal and practical consequences.
@@
Third, the predominance of analysis of fairness and discrimination in ML from the United States (US), lack of non-US ML examples (Le Quy et al., 2022), and limited legal scholarship translating these concepts across other jurisdictions, has inadvertently fostered a series of misconceptions on what would be unlawful algorithmic discrimination. Few papers have engaged with anti-discrimination laws outside the US (see e.g., Xenidis, 2020; Lattimore et al., 2020; Wachter et al., 2021b; Weerts et al., 2023; Kiesow Cortez and Maslej, 2023; Langenbucher, 2023), and fewer still in the United Kingdom (UK(see Sánchez-Monedero et al., 2020; Adams-Prassl et al., 2022; Kelly-Lyth, 2021). By avoiding the nuanced legal realities of other jurisdictions, models designed to comply with US laws may breach UK laws or those in comparable jurisdictions. English common law is either in force or is the dominant influence in 80 legal systems that govern approximately 2.8 billion people, not including the US (CIA, 2024). In particular, UK anti-discrimination law is very similar to numerous Commonwealth and common law jurisdictions, including Australia (AHRC, 2014), Canada (Canadian Parliament, 1985), India (Indian Parliament, 1950), New Zealand (New Zealand Parliament, 1993), South Africa (South African Parliament, 2000), and the pending bill in Bangladesh (Bangladesh Parliament, 2022). European Union (EU) law also has broadly the same discrimination law as it evolved in parallel during the UK’s membership (European Union, 2009).
@@
As new AI regulations emerge worldwide to prevent discriminatory practices, the need for a more jurisdictionally nuanced understanding of unlawful discrimination decision-making based on SML is imperative (European Parliament, 2024; Biden, 2023). Our paper addresses this gap by providing a rigorous analysis of UK anti-discrimination law, correcting certain mischaracterisations, and establishing a more accurate foundation for developing non-discriminatory ML in the UK and related jurisdictions.
@@
    We propose “conditional estimation parity” as a new, legally informed target to minimise the legal and practical effects of estimation error in SML models.
@@ 2.1. Automated Decision-Making
Let $x_{i}\in\mathbb{R}^{p}$ be a vector of observed attributes for individual $i$. A decision-maker must choose a decision $a\in\mathcal{A}$, where $\mathcal{A}$ is closed. We assume the decision-maker wants to decide based on a future outcome $y_{i}\in\mathcal{Y}$ for individual $i$. Further, we assume $\mathcal{Y}=\mathbb{N}$, which can be relaxed. Decision-making under uncertainty has long been studied in statistical decision theory (Savage, 1956; DeGroot, 1970; Berger, 1985; Parmigiani and Inoue, 2010). In a decision setting, let $u(y,a)$ be a utility function that summarises the utility for the decision-maker, where the optimal decision under uncertainty is
@@ 2.2. Algorithmic Fairness
Algorithmic fairness metrics generally measure prediction disparities across groups with different legally protected characteristics commonly identified in datasets, including gender and race (Hardt et al., 2016; Corbett-Davies et al., 2017; Narayanan, 2018; Verma and Rubin, 2018; Mulligan et al., 2019; Binns, 2018; Mehrabi et al., 2021; Mitchell et al., 2021). This research has resulted in several proposals, including statistical metrics to assess the fairness of individual predictive models (Verma and Rubin, 2018; Speicher et al., 2018; Carey and Wu, 2023; Caton and Haas, 2023), fairness for model auditing (Kim, 2017; Raji et al., 2020; Kasy and Abebe, 2021; Mökander, 2023; O’Neil et al., 2024), and fairness constraints on models (Corbett-Davies et al., 2017; Zafar et al., 2019; Heidari et al., 2019; Xu et al., 2020; Becker et al., 2023). We outline two core metrics that are relevant to our work – statistical parity and conditional statistical parity. To define these metrics in our notation, we separate $x_{i}$ into protected and legitimate features $x_{i}=(x_{pi},x_{li})$; we drop $i$ to simplify notation. Here, $x_{p}\in\mathcal{C}$ indicates protected attributes, with $\mathcal{C}$ being the set of different categories or groups.
Statistical parity, or demographic parity, is one of the central algorithmic fairness metrics (Corbett-Davies et al., 2017; Verma and Rubin, 2018; Narayanan, 2018; Mehrabi et al., 2021). For statistical parity to hold, it requires that
@@
such that the model predictions, in expectation over $x$, need to be the same for the different groups (Corbett-Davies et al., 2017; Verma and Rubin, 2018). Given that the decision function $d(\boldsymbol{\pi})$ is the same for the different groups, statistical parity results in equal decisions across those groups.
@@
so there should be no difference in model predictions or decisions between groups given by the protected attribute, conditional on legitimate features $x_{l}$ (Corbett-Davies et al., 2017; Verma and Rubin, 2018; Castelnovo et al., 2022).
Other related group comparison metrics include error parity, balanced classification rates, and equalised odds (Hardt et al., 2016; Corbett-Davies et al., 2017; Verma and Rubin, 2018; Narayanan, 2018; Mehrabi et al., 2021; Corbett-Davies et al., 2023). Individual approaches to parity have also considered whether otherwise identical individuals are treated differently if they have different protected attributes (Dwork et al., 2012; Kilbertus et al., 2018; Binns, 2020). Concepts from causal inference and counterfactual reasoning have been proposed to measure outcome consistency for individuals across protected groups (Kusner et al., 2017; Kilbertus et al., 2017; Russell et al., 2017; Zhang and Bareinboim, 2018; Chiappa, 2019; Wu et al., 2019; Nilforoshan et al., 2022; Alvarez and Ruggieri, 2023). It is not in the scope or aim of this paper to evaluate these numerous algorithmic fairness metrics.
@@ 2.3. Anti-Discrimination Law and Decision-Theoretic Approach
Identifying a discriminatory AI system has occurred in a variety of contexts, often without identifying whether it is unlawfully discriminatory. A non-lawyer may be surprised by several types of unfair behaviours that are not legally prohibited and some inconspicuous decisions that result in unlawful discrimination (Khaitan, 2015). As a term that has entered the common vernacular, despite its specific meanings in several contexts, it has led to the conflation of discrimination and unlawful discrimination. Importantly, anti-discrimination law only applies to a select group of duty-bearers (Khaitan, 2015). For example, an individual choosing not to be friends with people based on their sexuality or choosing not to marry someone based on their race is not unlawful (Emens, 2009; Khaitan, 2015). Despite certain actions appearing unreasonable or unfair, they are not always prohibited. Consider an algorithm that rejects a loan application because the applicant uses an Android phone rather than an iOS device (Aggarwal, 2021; Langenbucher, 2023). While it may show a correlation to the applicant’s income, one may agree that it is unfair because it does not reflect the individual likelihood of defaulting on a loan and instead penalises the individual for how much money they spent on a mobile device. However, under UK law, it would not be unlawful because spending habits, income, or even poverty, are not protected attributes (OHCHR, 2022). Therefore, to engage in a functional analysis of UK anti-discrimination laws in automated decision systems, it is helpful to frame its legal elements.
@@
In most algorithmic fairness literature, the primary, or often the sole, consideration is whether the ML outputs result in disparate predictions or decisions for a protected group (Hardt et al., 2016; Verma and Rubin, 2018; Narayanan, 2018). Such approach only considers aspects of the rule articulation function of anti-discrimination law. However, we take a wholistic, legally-informed approach:
@@
We formalise UK anti-discrimination law in a decision-theoretic framework that provides a systematic approach for the decision-maker to make optimal ML design choices under uncertainty and under legal constraints. Our decision-theoretic view, where actions map states to outcomes, each with associated utilities, enables a formal assessment of modelling choices under UK anti-discrimination law.
@@ 3. What types of discrimination does the law prohibit?
Unlawful and prohibited discrimination is defined by (1) protected contexts, (2) protected characteristics, and (3) prohibited conduct. Protected contexts are defined by the imposed duties on government, employers, landlords, providers of goods and services (United Kingdom Parliament, 2010). The protected characteristics under the Equality Act are age, disability, gender reassignment, marriage and civil partnership, pregnancy and maternity, race, religion or belief, sex and sexual orientation (United Kingdom Parliament, 2010, s 4). Two types of conduct are prohibited: direct discrimination and indirect discrimination.
@@ 3.1. Algorithmic Direct Discrimination
Less favourable treatment of an individual in a protected context based on one or multiple protected characteristics is unlawful direct discrimination (United Kingdom Parliament, 2010, s 13(1), 14). Where a model $\hat{p}(y|x)$ uses a protected attribute $x_{p}$, and there is a difference in predictions between the protected groups defined by $x_{p}$ that results in less favourable treatment for an individual in that group, this risk arises.
In US literature, the prevailing view is that the analogous type of direct discrimination, “disparate treatment”, will be challenging to prove in an algorithmic context (Barocas and Selbst, 2016). However, the dominance of US legal framing has created an assumption that direct discrimination is similarly “likely to be less important” (Wachter et al., 2021a),(see also Kelly-Lyth, 2021; Xenidis, 2020; Zuiderveen Borgesius, 2020). This assumption is incorrect. As Adams-Prassl et al. (2022) explain, “the scope of direct discrimination is significantly wider than that of disparate treatment.”
First, UK direct discrimination focuses on whether a protected characteristic is the reason for less favourable treatment (United Kingdom Supreme Court, 2017b); it does not take a formalistic view of whether $x_{p}$ is considered in the decision-making in the “input-focused disparate treatment” doctrine of the US (cf. Gillis, 2022). The formalistic approach in the US resulted in computer science literature incorrectly encouraging the removal of protected attributes when designing ML (Romei and Ruggieri, 2014; Johndrow and Lum, 2019; Grabowicz et al., 2022),(cf. Yang and Dobbie, 2020; Hellman, 2020). However, in the UK there is no general prohibition on the knowledge or consideration of a protected characteristic, and even if a model ignores $x_{p}$, in practice, direct discrimination can also be based on a criterion that is some “indissociable” proxy which has an “exact correspondence” to the protected characteristic $x_{p}$ (United Kingdom Supreme Court, 2017c, 2018). Formally, we can define an exact proxy as a feature $\tilde{x}_{p}$ with an exact correspondence with $x_{p}$. UK Courts have held exact proxies to include the criterion of statutory retirement age that differed between men and women as a proxy for sex (United Kingdom House of Lords, 1990), the criterion of marriage was historically indissociable from heterosexual orientation (United Kingdom Supreme Court, 2017a), or pregnancy as an exact proxy to the female sex (United Kingdom House of Lords, 1995; United Kingdom Employment Appeals Tribunal, 1996). This approach to direct proxy discrimination also applies under EU law (Weerts et al., 2023; European Court of Justice, 1990, 2008b, 2008a). If a model uses such features, it would have the effect of using an exact proxy $\tilde{x}_{p}$ that could be the basis for a direct discrimination claim. Therefore, removing $x_{p}$ will not avoid liability for unlawful direct discrimination.
@@
Further, from a technical perspective simply removing protected characteristics may reduce accuracy and utility (Zhang et al., 2017; Khani and Liang, 2021), and does not remove the risk of bias or disparity (Dwork et al., 2012; Lipton et al., 2018; Kleinberg et al., 2019). If the inclusion of $x_{p}$ or $\tilde{x}_{p}$ improves the model accuracy without resulting in less favourable treatment for protected individuals, it may avoid direct discrimination. There is an absence of any legal guidance in the UK on the use of protected attributes in automated decision-making. [^1] Although in the context of data processing, it may be lawful to use data on personal characteristics if necessary to identify or review equality of opportunity or treatment between groups of people, or to prevent or detect other unlawful acts (United Kingdom Parliament, 2018, Sch 1). Pending further legal guidance, it is important to carefully consider the effects of including $x_{p}$ or $\tilde{x}_{p}$.
@@
Second, intention is irrelevant to direct discrimination, “no hostile or malicious motive is required” (United Kingdom Supreme Court, 2017b; United Kingdom House of Lords, 1990; United Kingdom Supreme Court, 2009). Importantly, this diverges from US law and highlights that intention is immaterial to UK direct discrimination and to consideration of intentional proxy discrimination (cf. Prince and Schwarcz, 2020; Tschantz, 2022).